Status Report

NASA Space Station On-Orbit Status 30 January 2006

By SpaceRef Editor
January 30, 2006
Filed under , , ,
NASA Space Station On-Orbit Status 30 January 2006

  SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2006) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Underway: Week 17 for Expedition 12.

Onboard sleep cycle continues to be shifted a total of 7 hours to the right (wake up 8:00am EST, sleep at 11:30pm), to prepare the crew for this week s EVA-15 spacewalk (which begins at ~5:26pm EST).

The Science Officer continued his support of the PromISS-4 (Protein Crystal Growth Monitoring by Digital Holographic Microscope #4) experiment, today removing tape #16 from the MSG (Microgravity Science Glovebox) video system. Later, McArthur inserted tape #1 for PromISS-4.  [Tape exchanges vary due to record times throughout the 15 day running period of the experiment.]

McArthur and Tokarev continued preparations for the Orlan comm check today, the suited dry run tomorrow (1/31) and the EVA-15 next Friday (2/3).

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV


EVA preparations today included: (1) attaching the LCG (liquid cooling garment) restraint straps of the Orlan-Ms; (2) suit fit adjustment for height (to be repeated as necessary during the dry-run and under reduced airlock pressure); (3) leak checks and valve functionality tests on the suits, their BSS interface units and the hatch KVDs (pressure equalization valves, U.S.: PEVs) in the Pirs Docking Compartment (DC1) and Service Module Transfer Compartment (SM PkhO) from their EVA support panels (POV); (4) leak checking on the oxygen repressurization tanks (BK-3, primary & backup); (5) filling the DIDBs (disposable in-suit drink bags) and installing them in the suits, and (6) installing the Fresnel lens viewing aid in the helmets (at crew s discretion).

After setting up communications links via the suits radio telemetry units (BRTA-13 for Valery s Orlan #27, BRTA-12 for Bill s Orlan #25), the crew performed checks on Orlan and BSS telemetry, voice, and biomedical parameter transmission, testing the Russian BETA-08 ECG (electrocardiogram) lead cable belts, worn under the Orlan-M suits, for their function, using the Gamma-1M medical complex from the PKO medical exam panel. Testing of Orlan Korona communication links included VHF/voice and biomedical electrode belt and telemetry hookups via the BSS (later to be via the wireless in-suit radio telemetry system Tranzit-B) for vital signs and equipment monitoring.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

For the comm checks, the remote-commanded LKT1G3 commutator (switch) disconnected the Progress 20 telemetry from the BITS2-20 onboard telemetry system. After completion of the DC1 operations, FE Tokarev reset the communications system to its regular configuration.

The crew also checked out and installed the U.S. add-on hardware on the Orlan-M suits.  [Add-ons include the OTA consumables, right swing arm, a small trash bag, wire ties and safety tethers.]

The CDR started the charge/discharge cycle on the EMU batteries in the U.S. Airlock.

McArthur conducted routine maintenance on the new CSA-CP (Compound Specific Analyzer-Combustion Products) units, #1012 & #1017, currently in use as prime and backup units.  [As usual, the #1012 prime unit received a new battery, and both units were zero-calibrated. The backup unit was then returned to its location in the Node. After data take, the data logger was deactivated again.]

Later, McArthur performed the daily atmospheric status checks for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen Sensor).  [Over the last month, ground engineers have noticed a discrepancy between the SM GA (Gas Analyzer) ppCO2 reading and that of the MCA (Major Constituents Analyzer). In order to help assess the situation, TsUP-Moscow had the Soyuz GA powered on to get another sensor’s readings. McArthur used the CDMK (CO2 Monitoring Kit) for the same reasons. In addition to the ppCO2 issue described above, there appears to be also a trending apart of the MCA ppO2 value and the CSA-CP O2 sensor value. Therefore, starting on 1/26, the CDR also is taking daily CDMK and CSA-O2 readings to assist in evaluating these trends. Both MCC-H and TsUP are evaluating these issues. Despite the discrepancies, the atmosphere on ISS is considered in a safe configuration.]

FE Tokarev performed the daily routine maintenance of the SM’s environment control & life support system (SOZh), including its toilet system (ASU).

Both crewmembers completed their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive exerciser and VELO bike with bungee cord load trainer.  [Valery s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today: Day 1 of the first set).]

Later today, the FE will check the operation of the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).  [This daily monitoring/temp checking, carried on the Russian voluntary “time available” task list, will continue until 4/30.]

Working off his discretionary time available task list, Tokarev worked on the BIO-5 Rasteniya-2 (“Plants-2”) experiment, completing the daily status check and the regular periodic photographing of the experiment s progress for temporary computer storage and subsequent downlink to the ground. He also topped off the water tank.  [Rasteniya researches growth and development of plants under spaceflight conditions in the Lada-8 greenhouse. The regular maintenance of the experiment (each Monday, Wednesday, Friday and Sunday) involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, topping off the water tank if ~20-25% of the total amount (4 liters) remains, and photo/video recording. Once weekly, data from the Lada greenhouse control unit are recorded on floppy disk for weekly downlink via REGUL-Packet or the new BSR-TM at a suitable occasion.

Tonight at ~10:40pm EST, the station will maneuver from LVLH YVV (local vertical local horizontal/y-axis in velocity vector) to LVLH +XVV (+x-axis in velocity vector).

No CEO (Crew Earth Observation) photo targets uplinked today.

Events Ahead (all dates Eastern; tentative):

  • 01/31/06 — Orlan spacesuit dry run
  • 02/03/06 — Russian EVA-15
  • 02/11/06 — ISS Reboost Test (in MMOD avoidance mode)
  • 03/03/06 — Progress M-54/19P undocking & reentry
  • 03/30/06 — Soyuz TMA-8/12S launch (Exp. 13 + Marcos Pontes/Brazil)
  • 04/01/06 — Soyuz TMA-8/12S docking (DC1)
  • 04/24/06 — Progress M-56/21P launch
  • 04/26/06 — Progress M-56/21P docking
  • 06/19/06 — Progress M-55/20P undocking & reentry.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.